OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 can someone please help?? @dan815 @bakonloverk @mathstudent55 @ObeySoph▲ @PRAETORIAN.10 someone??? please???

OpenStudy (anonymous):

x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution these are the answer choices

OpenStudy (anonymous):

@ganeshie8 @lxl-rhea-lxl @AnimalAin @zepdrix

OpenStudy (anonymous):

@firstyoufall @jdoe0001 @KlOwNlOvE @Abmon98 @hba

zepdrix (zepdrix):

\[\Large\rm |2x + 6| − 4 = 20\]Add 4 to each side,\[\Large\rm |2x + 6|= 24\]When you see absolute bars, you can drop them by tossing a plus/minus sign into the mix.\[\Large\rm 2x + 6= \pm24\]

zepdrix (zepdrix):

So you now have two equations. One with a positive 24,\[\Large\rm 2x + 6= 24\]And another with a negative 24,\[\Large\rm 2x + 6= -24\]

zepdrix (zepdrix):

Do you understand how to solve for x in each case?

OpenStudy (anonymous):

yes I do. I first subtract six from -24 which is -30 and divide it by 2 which is -15. and the one with the positive I divide 6 from positive 24 which is 18 and divide it by 2 which is 9

OpenStudy (anonymous):

so is it -15 and 9??

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

can you help me with another one? I will definitely give you a medal afterwards

zepdrix (zepdrix):

lol sure c:

zepdrix (zepdrix):

Holding that medal hostage huh? >.< I see how it is lol

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day? x = −9, x = 21 x = −6, x = 18 x = 6, x = −18 no solution

zepdrix (zepdrix):

Same as the last one ya dingus! I would recommend isolating the absolute stuff before dropping the bars.

OpenStudy (anonymous):

I took this test already 10 times but I always only get 3/10 right

OpenStudy (anonymous):

ok well thank you

OpenStudy (anonymous):

for the answers I get 11 and -27??????

zepdrix (zepdrix):

\[\Large\rm 2|x − 6|+14=38\]Subtract 14 from each side,\[\Large\rm 2|x − 6|=24\]Divide by 2,\[\Large\rm |x − 6|=12\]Drop the jail bars,\[\Large\rm x-6=\pm12\]Hmm no those don't seem to be the answers we're looking for.

OpenStudy (anonymous):

18 and -6??

zepdrix (zepdrix):

yay good job \c:/

zepdrix (zepdrix):

Getting confused by some of those algebra steps maybe?

OpenStudy (anonymous):

weird. I still only got 5/10. but thank u so much. the ones u helped with I got right tho c:

zepdrix (zepdrix):

aw :c is it FLVS where you have to get at least 8 or something?

OpenStudy (anonymous):

yea it really sucks.